OpenStudy (anonymous):

The diameters of two semi-circles are 40 cm and 46 respectively. Find their total area.

OpenStudy (anonymous):

well remember the area of a circle formula is just: \[A = \pi r^2 \] and remember that radius is just half the diameter, since we have the diameter, we can plug it in to find the radius r = 1/2D r = 1/2 (40) = 20 \[A = \pi (20)^2 \]\[= \pi 400\] using this same method we can find the other circle: r = 1/2D r = 1/2 (46) = 23 \[A = \pi (23)^2 \] \[= \pi529 \]

OpenStudy (anonymous):

so how do i get the total area?

OpenStudy (anonymous):

oh wait its a semi circle, we have to halve the our results: \[A_{1}= 1/2(\pi400) = \pi 200\] \[A _{2}= 1/2(\pi529) = \pi264.5\]

OpenStudy (anonymous):

I'm not sure what they mean by total area. Do they want the area of each semi-circle or the combined area? if it is combined then add them together

OpenStudy (anonymous):

o alright thanks its combined area.